The electric field and the magnetic field in free space are given by




E=\frac{50}{\rho}\cos(10^{6}t+\beta z)a_{\phi} V/mE=



ρ



50





cos(10



6



t+βz)a



ϕ





V/m




H=\frac{H_{o}}{\rho}\cos(10^{6}t+\beta z)a_{\rho} A/mH=



ρ



H



o









cos(10



6



t+βz)a



ρ





A/m




Express these in phasor form and determine the constants H_{o}H



o





and \betaβ such that the fields satisfy Maxwell’s equations.

1
Expert's answer
2021-11-21T17:24:53-0500

"E=\frac{50}{\rho}\cos(10^{6}t+\beta z)a_{\phi} \frac Vm,"

"H=\frac{H_o}{\rho}\cos(10^{6}t+\beta z)a_{\rho}~ \frac Am,"

"H_o=50,~\beta=1."